# The intrinsic subgroup of an elliptic curve and Mazur’s torsion theorem

### v1

@@ Abstract.
We define and study a biadditive symmetric (not necessarily perfect) pairing on the torsion part $\mathrm{Pic}(X)_{{\operatorname{tors}}}$ of the Picard group of a smooth projective curve $X$ over a field $k$ with values in $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$. We call its kernel the intrinsic subgroup of $X$. It turns out that some information on the reduction type of $X$ can be read off from the intrinsic subgroup. Mazur’s torsion theorem says that there are exactly 15 isomorphism classes of abelian groups that appear as the rational torsion points of an elliptic curve $X$ over $\mathbb{Q}$ (identified with $\mathrm{Pic}(X)_{\operatorname{tors}}$). We refine this result by determining which subgroups of those 15 groups appear as the intrinsic subgroups.
@@ 5.1. Modular curve $X_{1}(N)_{M}^{\pm}$
(2) We first note that the cases $(N,M)\in\{(8,8),(9,9),(10,10),(12,4),(12,6),(12,12)\}$ are reduced to the case $(N,M^{\prime})$ for some smaller $M^{\prime}|N$ by Lemma 5.4 (3). Next, if $\epsilon=+$ or $M$ is odd, we can apply Remark 5.2 (1) and Corollary 5.7 to deduce $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})=\emptyset$. This covers the cases $(M,N)\in\{(7,7),(10,5),(12,3)\}$ with $\epsilon\in\{+,-\}$, and $(M,N)\in\{(6,6),(8,4),(10,2),(12,2)\}$ with $\epsilon=$-.
@@ 6.2. Proof of existence
Let $(E,P,Q)\in Y_{1}^{0}(4,2)(\mathbb{Q})$, and suppose first $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ is non-trivial. Then we must have $E(\mathbb{Q})_{\operatorname{tors}}\cong A$ (because we have shown $E(\mathbb{Q})_{\operatorname{tors}}\cong\mathbb{Z}/8\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ implies $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$), and then $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|=2$ (because we have shown $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|\not=4,8$). Thus we have
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle 2P\rangle\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$,
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle Q\rangle\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$.
Since the components of $Y_{1}^{0}(N)_{(2,1,1)}^{\pm}$ and $Y_{1}^{0}(N)_{(1,2,2)}^{\pm}$ are isomorphic to $\mathbb{P}^{1}$ by Proposition 5.15, we conclude by Proposition 6.2 that there are infinitely many such $(E,P,Q)$ in both cases. Finally, we have $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$ precisely when neither of the above conditions are met and $(E,P,Q)\not\in\operatorname{Im}(Y_{1}^{0}(8,2)(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$. Again we conclude by Proposition 6.2 that there are infinitely many such $(E,P,Q)$.
@@
$J:=$ $\operatorname{Im}(Y_{1}(2,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{1}:=$ $\operatorname{Im}(Y_{1}(4,2)(\mathbb{Q})\sqcup Y_{1}(6,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{2}:=$ $\operatorname{Im}(Y_{1}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\sqcup Y_{1}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{3}:=$ $\{j(E)\in Y_{1}(1)(\mathbb{Q})\mid(E,P,Q)\in Y_{1}(2,2)(\mathbb{Q}),\ E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\not=0\}.$

### v2

@@ Abstract.
We define and study a biadditive symmetric (not necessarily perfect) pairing on the torsion part $\mathrm{Pic}(X)_{\mathrm{tors}}$ of the Picard group of a smooth projective curve $X$ over a field $k$ with values in $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$. We call its kernel the intrinsic subgroup of $X$. It turns out that some information on the reduction type of $X$ can be read off from the intrinsic subgroup. Mazur’s torsion theorem says that there are exactly 15 isomorphism classes of abelian groups that appear as the rational torsion points of an elliptic curve $X$ over $\mathbb{Q}$ (identified with $\mathrm{Pic}(X)_{\mathrm{Tor}}$). We refine this result by determining which subgroups of those 15 groups appear as the intrinsic subgroups.
@@ 5.1. Modular curve $X_{1}(N)_{M}^{\pm}$
(2) We first note that the cases $(N,M)\in\{(8,8),(9,9),(10,10),(12,4),(12,6),(12,12)\}$ are reduced to the case $(N,M^{\prime})$ for some smaller $M^{\prime}|N$ by Lemma 5.4 (3). Next, if $\epsilon=+$ or $M$ is odd, we can apply Remark 5.2 (1) and Corollary 5.7 to deduce $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})=\emptyset$. This covers the cases $(N,M)\in\{(7,7),(10,5),(12,3)\}$ with $\epsilon\in\{+,-\}$, and $(N,M)\in\{(6,6),(8,4),(10,2),(12,2)\}$ with $\epsilon=$-.
@@ 6.2. Proof of existence
Let us introduce three subsets of $Y_{1}^{0}(4,2)(\mathbb{Q})$:
$J_{1}:=$ $\operatorname{Im}(Y_{1}^{0}(8,2)(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q})),$
$J_{2}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)^{\pm}_{(2,1,1)}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q})),$
$J_{3}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)^{\pm}_{(1,2,2)}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q})).$
Note that $J_{1}\cap J_{3}=\emptyset$. Indeed, since $Y_{1}^{0}(8,2)_{(1,2,1)}^{\pm}(\mathbb{Q})=\emptyset$ by Proposition 5.15, we have $\langle Q,Q\rangle\not=0$ for any $(E,P^{\prime},Q)\in Y_{1}^{0}(8,2)^{\pm}(\mathbb{Q})$, which implies the claim.
For $(E,P,Q)\in Y_{1}^{0}(4,2)(\mathbb{Q})$, we have $E(\mathbb{Q})_{\operatorname{tors}}=\langle P,Q\rangle\Leftrightarrow(E,P,Q)\not\in J_{1}$. Suppose this holds. Then we know that $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|\leq 2$ (by the non-existence part). Therefore we have the equivalences
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle 2P\rangle\Leftrightarrow(E,P,Q)\in J_{2}$,
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle Q\rangle\Leftrightarrow(E,P,Q)\in J_{3}$,
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0\Leftrightarrow(E,P,Q)\not\in J_{2}\cup J_{3}$.
We conclude the existence part of Theorem 1.1 in this case by Proposition 6.2.
@@
$J:=$ $\operatorname{Im}(Y_{1}^{0}(2,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{1}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)(\mathbb{Q})\sqcup Y_{1}^{0}(6,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{2}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\sqcup Y_{1}^{0}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{3}:=$ $\{j(E)\in Y_{1}(1)(\mathbb{Q})\mid(E,P,Q)\in Y_{1}^{0}(2,2)(\mathbb{Q}),\ E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\not=0\}.$